OpenStudy (anonymous):

Horizontal asymptote help! I got no horizontal asymptotes for both, but I think that's wrong. State the horizontal asymptote of the rational function. f(x) = (x^2 + 6x - 8) / (x - 8) and f(x) = (x^2 + 3x - 2) / (x - 2)

OpenStudy (anonymous):

the first: y=x+8 i'm doing the second, do u wonna the passages?

OpenStudy (anonymous):

sorry y=x+6

OpenStudy (anonymous):

So the horizontal asymptote really isn't zero, right? I got zero, but I have a feeling it's wrong.

OpenStudy (anonymous):

just a second, i'm checking out the calculus ;)

OpenStudy (anonymous):

Ok! thx

OpenStudy (anonymous):

Ok the first is : y=x+14 (for sure xD)

OpenStudy (anonymous):

Ok great

OpenStudy (anonymous):

Could you show me how you got that answer, if you have the time :)

OpenStudy (anonymous):

I should have got the time, butt someone enjoyed in suspending me because I didn't give you the explanation (actually I was typing it) :)

OpenStudy (anonymous):

\[m=limx-->\inf (fx/x)\] and that equals one

OpenStudy (anonymous):

Then, \[q=limx-->\inf(fx-mx)\] and that equals 14 So y=x+14 If you haven't got with the limits tell me ;)

OpenStudy (anonymous):

The other one is the same method of course, it gives me y=x+5 The same, tell me if you have problem with calculus ;) ****PLEASE DO NOT SUSPEND ME I GAVE THE FOURMULA AND EXPLANATION ABOVE****

OpenStudy (anonymous):

Thank you for your time and sorry for the trouble!

OpenStudy (anonymous):

Don't worry it's not your fault :)